# Deep reinforcement learning-based active flow control of an elliptical cylinder: transitioning from an elliptical cylinder to a circular cylinder and a flat plate

### v3

@@ Abstract
This study investigates the effectiveness of active flow control (AFC) technology supported by deep reinforcement learning (DRL) applied to flows around an elliptical cylinder at $Re=100$. We vary the aspect ratio ($Ar$) of the elliptical cylinder from an elliptical cylinder ($Ar=2.0$) to a circular cylinder ($Ar=1.0$), and ultimately to a flat plate ($Ar=0$). We utilize the Proximal Policy Optimization (PPO) algorithm to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of a cylinder to achieve reduction in drag, minimization of lift, and suppression of vortex shedding. The robustness and adaptability of DRL-based control techniques are examined across different geometric configurations. Our research findings indicate that, for elliptical cylinders with $Ar$ between 1.75 and 0.75, the reduction in drag coefficient ranges from 0.9% to 15.7%, and the reduction in lift coefficient ranges from 95.2% to 99.7%. The DRL-based control strategy not only significantly reduces lift and drag, but also completely suppresses vortex shedding while using less than 1% of external excitation energy, demonstrating its efficiency and energy-saving capabilities. Additionally, for $Ar$ from 0.5 to 0, the reduction in drag coefficient ranges from 26.9% to 43.6%, and the reduction in lift coefficient from 50.2% to 68.0%. This reflects the control strategy’s significant reduction in both drag and lift coefficients, while also alleviating vortex shedding. It is expected underscore the adaptability and potential of DRL-based AFC in controlling complex fluid dynamics across diverse geometric configurations.
@@ I INTRODUCTION
The development of active flow control (AFC) technology has progressed from theoretical exploration to practical application, gradually becoming an important branch within the domains of fluid mechanics and engineering.Collis and Joslin (2004); Jahanmiri (2010); Awada, Younes, and Ilinca (2021) This technology actively intervenes the flow field to improve flow characteristics, such as reducing drag, controlling vortex shedding, enhancing mixing efficiency, or preventing flow separation, thereby enhancing system performance and efficiency.Collis and Joslin (2004); Wang et al. (2018) In its early stages, AFC research primarily focused on theoretical analysis and simple experiments to explore the influence of external inputs, such as jets, vibrations, or magnetic fields, on fluid flow dynamics.Jahanmiri (2010); Wang et al. (2018); Awada, Younes, and Ilinca (2021) With the advancement of computational fluid dynamics (CFD) and increased computational power, researchers have been able to delve deeper into complex flow responses through numerical simulations and develop more effective control strategies.Aram et al. (2018); Jain, Yeo, and Chopra (2010); Li et al. (2022) However, as fluid problems are typically characterized by a high degree of nonlinearity and multiscale nature, involving high-dimensional state spaces and complex system dynamicsPozorski and Wacławczyk (2020); Reichstein et al. (2019), many engineering problems are often intractable using existing models. Sophisticated control strategies that require consideration of a wide range of spatiotemporal scales are indispensable when addressing phenomena such as unstable flows, turbulence, and multiphase flows.Subramaniam (2020)
@@
Recent advances in machine learning and deep learning technologies have revolutionized the field of artificial fluid control Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Ren, Hu, and Tang (2020). Deep Reinforcement Learning (DRL) combines deep learning and reinforcement learning to optimize decision-making through interactions with the environment Arulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2020). DRL excels at extracting complex features from large datasets and making optimal decisions based on these features Janiesch, Zschech, and Heinrich (2021); LeCun, Bengio, and Hinton (2015). By using deep neural networks, DRL is capable of handling high-dimensional state spaces and enables machines to tackle previously intractable decision-making challenges Arulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2020). DRL’s nonlinear modeling capabilities and adaptive learning mechanisms allow it to learn complex strategies directly from raw sensory data without manual feature engineering Dargazany (2021). It continuously improves its strategies through interactions with the environment Wang et al. (2020), making it advantageous for dynamic decision-making problems François-Lavet et al. (2018); Wang et al. (2020). DRL has found applications in various fields including gaming, autonomous vehicles, robotics, natural language processing, and computer vision. In particular, it has achieved distinguished success, such as AlphaGo’s victory over world champions in gaming Granter, Beck, and Papke (2017), route planning and decision-making optimization in autonomous driving Sallab et al. (2017), and complex physical tasks learning in robotics Zhao, Queralta, and Westerlund (2020).
@@
With progressive advancement in the DRL technology, there has been a significant increase in the application of DRL in the field of fluid dynamics as well.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al. (2021); Razdan and Shah (2022) DRL methods are highly suited to handle the high-dimensional and complex state spaces intrinsic to fluid dynamics.Xie et al. (2023) Through the function approximation capabilities of deep neural networks, DRL is adept at processing fluid systems and discerning flow patterns and structures.Zhao et al. (2024) DRL models, through their interaction with the environment and trial-and-error processes, can independently learn and refine strategies to meet diverse fluid conditions.Husen, Chaudary, and Ahmad (2022); Collis and Joslin (2004); Jahanmiri (2010) For instance, DRL has been applied to shape optimization of fluid systems and enabled automatic discovery and application of effective shape adjustments to achieve optimization goals by intelligently exploring the design space and learning optimal strategies.Bhola et al. (2023) Moreover, the long-term reward mechanism intrinsic to DRL allows for the consideration of future impacts when formulating strategies. This unique feature has allowed DRL to shine in many AFC problems where DRL’s focus on long-term rewards enables control models to balance short-term gains with future outcomes, thereby better addressing control and optimization challenges within fluid mechanics.Vinuesa et al. (2022); Vignon, Rabault, and Vinuesa (2023); Viquerat et al. (2022); Vinuesa (2024)
@@
In literature, DRL has attracted considerable interest for flow control around blunt bodies, as flow around blunt body tends to involve strong non-linearities such as flow separation and recirculation, traditional control methods often struggle in achieving control objectives whereas DRL models have shown promising results. Pioneered by Rabault et al. Rabault et al. (2019), various researchers such as Tang et al. Tang et al. (2020), Heess et al. Heess et al. (2017), Ren, Rabault, and Tang Ren, Rabault, and Tang (2021), and Wang and Xu Wang and Xu (2024a) applied the Proximal Policy Optimization (PPO) algorithm to AFC in the flow around a circular cylinder. Consistently, they achieved an approximate 8% reduction in drag in their validation experiments. Furthermore, Li and ZhangLi and Zhang (2022) ,Wang et al.Wang et al. (2022a), Wang and XuWang and Xu (2024b) and Wang et al.Wang et al. (2022b)He et al. (2023) respectively conducted flow control studies on confined circular cylinder, square cylinder and airfoil using DRL. In addition, Fan et al.Fan et al. (2020) initiated the application of DRL to AFC experiments, optimizing power efficiency by adjusting the rotational speed of auxiliary circular cylinders positioned behind another circular cylinder. Ren, Wang, and TangRen, Wang, and Tang (2021) proposed an AFC strategy to conceal the hydrodynamic characteristics of circular cylinders, such as strong shear forces and periodic shedding vortices. Studies on flow around constrained circular cylinders under different blockage ratios have also received extensive attention in the field Rabault et al. (2019); Tang et al. (2020); Li and Zhang (2022); Wang et al. (2022a); Wang and Xu (2024a).
Although active flow control strategies for both circular and square cylinders are pivotal in various engineering applications for managing aerodynamic and hydrodynamic forcesVignon, Rabault, and Vinuesa (2023); Rabault et al. (2020); Garnier et al. (2021); Razdan and Shah (2022), their extension to more intricate, non-ideal geometries poses significant challenges. The effectiveness of DRL strategies in such contexts remains largely unexplored, marked by uncertainties regarding their capacity to generate physically reasonable and adaptable control measures. Current literature indicates a distinct lack of comprehensive studies on the variability of DRL-based flow control strategies across different geometries, such as from streamlined bodies to various bluff bodies. Based on the studies conducted by Jia and Xu Wang and Xu (2024a, b), we focus on the development of the DRL technqiue for active flow control around elliptical cylinders, with the objective of developing effective control strategies for drag reduction, flow stability enhancement, and overall performance improvement, addressing the research gap in applying DRL to different geometric shapes of elliptical cylinders. The structure of this paper is organized as follows: Sec. II introduces the problem description and provides a comprehensive procedure for implementing DRL. Sec. III presents the main results and analysis. Finally, Sec. IV concludes the paper by summarizing the key findings and discussing their implications.
@@ II.2 Numerical simulation
We utilize the open-source CFD package OpenFOAM as the incompressible flow solver, following the methodology described by Jasak et al.Jasak (2009); Jasak et al. (2007). This approach is widely employed due to its ability to provide reliable numerical algorithms for solving the Navier-Stokes equations. The solver implemented in OpenFOAM uses the finite volume method to discretize the computational domain into a mesh consisting of control volumes. To ensure numerical stability, a time step size of $\Delta t=0.0005$ is selected in the simulation.
@@ II.4 DRL-enhanced active flow control
Our AFC framework utilizes DRL algorithms, consisting of the agent and the environment. The agent, designed for continuous action control, employs algorithms like PPO or SAC. The environment chosen for interaction with the agent is the CFD simulation of the flow around an elliptical cylinder. The custom environment class, as presented by Rabault et al.Rabault et al. (2020), provides the main structure, and further details can be explored in their work. Figure 4 illustrates the fundamental framework of DRL, including the computational process of a single training episode and the utilized formulas.
@@
The control action $a_{t}$ involves blowing and suction of two synthetic jets on the elliptical cylinder. The flow rates of the upper and lower synthetic jets are constrained such that their sum is zero, i.e., $Q_{\text{lower}}+Q_{\text{upper}}=0$. The RL agent controls the velocity of the upper synthetic jet, while the lower synthetic jet has the same velocity but in the opposite direction. To ensure energy efficiency in AFC, the velocity of the synthetic jets is limited to less than 2% of the inlet mean velocity, preventing the agent from utilizing excessively high jet velocities for drag reduction. Each control action, known as a ‘viscous action’, has a duration of 50 time steps. To ensure smooth transitions between consecutive actions, a smoothing function $S$ is introduced:
@@
where $C_{D,0}$ represents the baseline drag coefficient, serving as a reference point for drag reduction. $(C_{D})_{T_{i}}$ denotes the drag coefficient at time step $T_{i}$, with the objective of minimizing this value relative to the baseline. $(C_{L})_{T_{i}}$ signifies the lift coefficient at time step $T_{i}$, whose absolute value is penalized to mitigate lift forces that may destabilize the flow around the bluff body. $\omega$ is a weighting factor that quantifies the trade-off between minimizing drag and controlling lift fluctuations.
The time granularity parameter significantly affects the agent’s performance and the complexity of the learning task. Precise adjustment to this hyperparameter, based on prior research and ongoing strategy refinement, is essential for achieving optimized control outcomes. In particular, each action in our analysis is applied for a duration of 0.025 seconds, which corresponds to 25 numerical simulation time steps. Each episode consists of 100 control actions, resulting in a maximum total time for an episode of $T_{\max}=T_{100}=2.5$ nondimensional time units. The shedding frequency of vortices around the elliptical cylinder in this study ranges from 0.3 to 0.4, which corresponds to 6.25 to 8.33 vortex shedding cycles per episode. The length of episodes is intentionally chosen to encompass multiple shedding cycles, which facilitates effective learning of the control algorithm by allowing the agent to observe and adapt to the flow dynamics over several cycles.
@@
The DRL training and numerical simulation are conducted on a high-performance computing (HPC) system equipped with an Intel® Xeon® Platinum 8358 CPU, operating at 2.60 GHz, and comprising a total of 64 cores evenly distributed across two sockets, each containing 32 cores. Specifically, the numerical simulation is performed using version 8 of the open-source software platform OpenFOAM® version 8 Jasak et al. (2007). The deployment of DRL algorithms is based on theTensorforce library Schaarschmidt et al. (2018), while the definitions of artificial neural networks (ANNs) and gradient descent algorithms are facilitated using the TensorFlow open-source library Abadi et al. (2016). For the DRL implementation, the agent is loaded using the Tensorforce platform with the PPO algorithm, and the environment is constructed using the Gym interface Brockman et al. (2016) . Fig. 4 illustrates the use of CFD calculations as the environment for DRL training and shows the effects of using 60 parallel CFD environments to accelerate the training process. This parallel strategy is inspired by the studies of Rabault and KuhnleRabault and Kuhnle (2019a) and Wang et al.Wang et al. (2022a), and also incorporateds insights from the work of Wang et al.Wang et al. (2022a) For a more comprehensive understanding, readers are encouraged to consult their works and the references therein.

### v4

@@ Abstract
We study the adaptability of deep reinforcement learning (DRL)-based active flow control (AFC) technology for bluff body flows with complex geometries. It is extended from a cylinder with $Ar=1$ to a flat elliptical cylinder with $Ar=2$, slender elliptical cylinders with $Ar$ less than 1, and a flat plate with $Ar=0$. The robustness and adaptability of DRL-based control technology will be assessed under varying levels of flow instability around bluff bodies. We utilize the Proximal Policy Optimization (PPO) algorithm to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of a cylinder to achieve reduction in drag, minimization of lift, and suppression of vortex shedding. Our research findings indicate that, for elliptical cylinders with $Ar$ between 1.75 and 0.75, the reduction in drag coefficient ranges from 0.9% to 15.7%, and the reduction in lift coefficient ranges from 95.2% to 99.7%. The DRL-based control strategy not only significantly reduces lift and drag, but also completely suppresses vortex shedding while using less than 1% of external excitation energy, demonstrating its efficiency and energy-saving capabilities. Additionally, for $Ar$ from 0.5 to 0, the reduction in drag coefficient ranges from 26.9% to 43.6%, and the reduction in lift coefficient from 50.2% to 68.0%. This reflects the control strategy’s significant reduction in both drag and lift coefficients, while also alleviating vortex shedding. The interaction and nonlinear development of vortices in the wake of elliptical cylinders lead to complex flow instability, and DRL-based AFC technology shows adaptability and potential in addressing flow control problems for this type of bluff body flow.
@@ I INTRODUCTION
The development of active flow control (AFC) technology progresses from theoretical exploration to practical application, gradually becoming an important branch within the domains of fluid mechanics and engineering.Collis and Joslin (2004); Jahanmiri (2010); Awada, Younes, and Ilinca (2021) This technology actively intervenes the flow field to improve flow characteristics, such as reducing drag, controlling vortex shedding, enhancing mixing efficiency, or preventing flow separation, thereby enhancing system performance and efficiency.Collis and Joslin (2004); Wang et al. (2018) In its early stages, AFC research primarily focused on theoretical analysis and simple experiments to explore the influence of external inputs, such as jets, vibrations, or magnetic fields, on fluid flow dynamics.Jahanmiri (2010); Wang et al. (2018); Awada, Younes, and Ilinca (2021) With the advancement of computational fluid dynamics (CFD) and increased computational power, researchers are able to delve deeper into complex flow responses through numerical simulations and develop more effective control strategies.Aram et al. (2018); Jain, Yeo, and Chopra (2010); Li et al. (2022) However, due to the high degree of nonlinearity and multiscale nature of fluid problems, which involve high-dimensional state spaces and complex system dynamics, many engineering problems are often intractable using existing models.Pozorski and Wacławczyk (2020); Reichstein et al. (2019); Vinuesa (2024) Sophisticated control strategies that require consideration of a wide range of spatiotemporal scales are indispensable when addressing phenomena such as unstable flowsVinuesa (2024), turbulenceRen, Rabault, and Tang (2021), and multiphase flows.Subramaniam (2020)
@@
State-of-the-art machine learning technologies profoundly impact the field of active fluid control.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Ren, Hu, and Tang (2020) Deep Reinforcement Learning (DRL) combines deep learning and reinforcement learning to optimize decision-making through interactions with the environment.Arulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2020) DRL excels at extracting complex features from large datasets and making optimal decisions based on these features.Janiesch, Zschech, and Heinrich (2021); LeCun, Bengio, and Hinton (2015) By using deep neural networks, DRL handles high-dimensional state spaces and enables machines to tackle previously intractable decision-making challenges.Arulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2020) DRL’s nonlinear modeling capabilities and adaptive learning mechanisms allow it to learn complex strategies directly from raw sensory data without manual feature engineering.Dargazany (2021) It continuously improves its strategies through interactions with the environment, making it advantageous for dynamic decision-making problems.François-Lavet et al. (2018); Wang et al. (2020) DRL finds applications in various fields including gaming, autonomous vehicles, robotics, natural language processing, and computer vision. In particular, it achieves distinguished success, such as AlphaGo’s victory over world champions in gaming, route planning and decision-making optimization in autonomous driving, and learning complex physical tasks in robotics.Granter, Beck, and Papke (2017); Zhao, Queralta, and Westerlund (2020); Sallab et al. (2017)
@@
With progressive advancements in DRL technology, there is a significant increase in the application of DRL in the field of fluid dynamics as well.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al. (2021); Razdan and Shah (2022) DRL methods are highly suited to handle the high-dimensional and complex state spaces intrinsic to fluid dynamics.Xie et al. (2023) Through the function approximation capabilities of deep neural networks, DRL is adept at processing fluid systems and discerning flow patterns and structures.Zhao et al. (2024) DRL models, through their interaction with the environment and trial-and-error processes, independently learn and refine strategies to meet diverse fluid conditions.Husen, Chaudary, and Ahmad (2022); Collis and Joslin (2004); Jahanmiri (2010) For instance, DRL is applied to shape optimization of fluid systems, enabling the automatic discovery and application of effective shape adjustments to achieve optimization goals by intelligently exploring the design space and learning optimal strategies.Bhola et al. (2023) Moreover, the long-term reward mechanism intrinsic to DRL allows for the consideration of future impacts when formulating strategies. This unique feature allows DRL to excel in many AFC problems, where DRL’s focus on long-term rewards enables control models to balance short-term gains with future outcomes, thereby better addressing control and optimization challenges within fluid mechanics.Vinuesa et al. (2022); Vignon, Rabault, and Vinuesa (2023); Viquerat et al. (2022); Vinuesa (2024)
@@
In the literature, DRL attracts considerable interest for flow control around bluff bodies, as flow around these bodies tends to involve strong nonlinearities such as flow separation and recirculation. Traditional control methods often struggle to achieve control objectives, whereas DRL models show promising results. Pioneered by Rabault et al. Rabault et al. (2019), various researchers such as Tang et al. Tang et al. (2020), Heess et al. Heess et al. (2017), Ren, Rabault, and Tang Ren, Rabault, and Tang (2021), and Wang and Xu Wang and Xu (2024a) apply the Proximal Policy Optimization (PPO) algorithm to AFC in the flow around a circular cylinder. Consistently, they achieve an approximate 8% reduction in drag in their validation experiments. Furthermore, Li and ZhangLi and Zhang (2022), Wang et al.Wang et al. (2022a), Wang and XuWang and Xu (2024b), Wang et al.Wang et al. (2022b), and He et al.He et al. (2023) respectively conduct flow control studies on confined circular cylinders, square cylinders, and airfoils using DRL. In addition, Fan et al.Fan et al. (2020) initiate the application of DRL to AFC experiments, optimizing power efficiency by adjusting the rotational speed of auxiliary circular cylinders positioned behind another circular cylinder. Ren, Wang, and TangRen, Wang, and Tang (2021) propose an AFC strategy to conceal the hydrodynamic characteristics of circular cylinders, such as strong shear forces and periodic shedding vortices. Studies on flow around constrained circular cylinders under different blockage ratios also receive extensive attention in the field.Rabault et al. (2019); Tang et al. (2020); Li and Zhang (2022); Wang et al. (2022a); Wang and Xu (2024a) The application of DRL-based AFC technology in bluff body flows is summarized in Tab. 1. DRL-based AFC control technology is successfully applied to square, circular, and airfoil flows. Additionally, multiple studies demonstrate that DRL-based flow control strategies exhibit robustness across different Reynolds numbers and achieve favorable control outcomes in fluid mechanics experiments. This indicates that DRL-based AFC technology provides effective control strategies in challenging high-dimensional and nonlinear scenarios.
Table 1: Application of AFC technology based on DRL in the flow around bluff body.
Re Bluff Body Control measures DRL algorithm Drag Reduction Reference
100 Cylinder Synthetic Jets PPO 8% Rabault et al.Rabault et al. (2019)
100 Cylinder Synthetic Jets PPO 8% Wang and XuWang and Xu (2024a)
100 Cylinder Synthetic Jets PPO 8% Wang et al.Wang et al. (2022a)
120 Cylinder Synthetic Jets S-PPO-CMA 18.4% Paris, Beneddine, and DandoisParis, Beneddine, and Dandois (2021)
100-300 Cylinder Synthetic Jets PPO 6%-24% He et al.He et al. (2023)
100-400 Cylinder Synthetic Jets PPO 6%-39% Tang et al.Tang et al. (2020)
1,000 Cylinder Synthetic Jets PPO 30% Ren, Rabault, and TangRen, Rabault, and Tang (2021)
10,160 Cylinder Two fast-rotating smaller cylinders TD3 30% Fan et al.Fan et al. (2020)
100 Square Synthetic Jets SAC 14% Wang et al.Wang et al. (2022a)
100 Square Synthetic Jets SAC 14% Chen et al.Chen et al. (2023)
100 Square Synthetic Jets SAC and TQC 17% Xia et al.Xia et al. (2024)
100-400 Square Synthetic Jets SAC 14%-47% Wang and XuWang and Xu (2024b)
100-500 Square Synthetic Jets SAC 14%-51% Jia and XuJia and Xu (2024)
500-2,000 Square Synthetic Jets SAC 44%-61% Yan et al.Yan et al. (2023)
The application of AFC strategies based on DRL is widely validated across various bluff body flow scenarios. Among these, research on confined circular cylinder flow under different blockage ratios is the most thorough. In contrast, studies related to elliptical cylinder flow are exceedingly rare in this domain. The aspect ratio of an elliptical cylinder significantly affects flow stability and vortex structure development, with the slender shape’s structural effects directly increasing the difficulty of flow control. While active flow control strategies for circular and square cylinders are crucial for managing aerodynamic and hydrodynamic performance in various engineering applications, extending these strategies to more complex, non-ideal geometries remains a major challenge.Vignon, Rabault, and Vinuesa (2023); Rabault et al. (2020) Under these circumstances, the effectiveness of DRL strategies is largely unexplored, particularly regarding whether they can produce physically reasonable and adaptable control measures. Current literature indicates a lack of comprehensive studies on the variability of DRL-based flow control strategies across different geometries, such as from streamlined bodies to various bluff bodies.Garnier et al. (2021); Razdan and Shah (2022) Therefore, this study preliminarily investigates whether it is possible to effectively reduce drag and potentially suppress vortex shedding behind an elliptical cylinder at $Re=100$. This lower Reynolds number allows for a detailed examination of the fundamental principles and mechanisms involved in the control strategy, enabling a rigorous evaluation without the additional complexities introduced by turbulence. The experiences and insights gained from this work are crucial for applying DRL to more practical engineering problems in fluid dynamics.
The structure of this paper is organized as follows: sec. II details the problem addressed in this study and describes the numerical simulation methods and deep reinforcement learning techniques used. Additionally, it presents the framework for solving the AFC problem using DRL, providing a thorough description of the fundamental elements involved. Sec. III presents the main results and analysis. In sec. III.1, the baseline flow for elliptical cylinders with different $Ar$ is described, and probe positions and distributions are designed using physical information from the flow field. Sec. III.2 provides a detailed description of the controlled process for an elliptical cylinder with $Ar=0.75$ during DRL training. We also illustrate the impact of synthetic jet actions, whether blowing or suction, on the streamlines around the elliptical cylinder using detailed streamline. Sec. III.3 systematically tests the flow control performance of elliptical cylinders with $Ar$ ranging from 0 to 2 using DRL-based AFC, verifying the robustness of DRL-based AFC technology for elliptical cylinders with different geometric shapes. Sec. III.4 analyzes the controlled flow to explore the physical mechanisms of control. Finally, sec. IV concludes the paper by summarizing the key findings and discussing their implications.
@@ II.2 Numerical simulation
We utilize the open-source CFD package OpenFOAM as the incompressible flow solver, following the methodology described by Jasak et al.Jasak (2009); Jasak et al. (2007). This approach is widely employed due to its ability to provide reliable numerical algorithms for solving the Navier-Stokes equations. The solver implemented in OpenFOAM uses the finite volume method to discretize the computational domain into a mesh consisting of control volumes. We choose the pisoFoam solver based on OpenFOAM, which is a transient solver for incompressible flow using the PISO algorithm. PISO algorithm is to iteratively correct the pressure and velocity fields to satisfy the incompressibility condition at each time step. The time and spatial derivatives are discretized using the Crank-Nicolson scheme and the second-order central difference scheme, respectively. A non-dimensional time step of $\Delta t=0.0005$ is selected. The CFL number, defined as $\left(\frac{\Delta t|U|}{\Delta x}\right)$—where $|U|$ represents the magnitude of the velocity vector per grid unit and $\Delta x$ is the grid length in the direction of the velocity—is less than 1, meeting the accuracy requirements.
@@ II.4 DRL-enhanced active flow control
Our AFC framework utilizes DRL algorithms, consisting of the agent and the environment. The agent, designed for continuous action control, employs algorithms like PPO or SAC. The environment chosen for interaction with the agent is the CFD simulation of the flow around an elliptical cylinder. The custom environment class, as presented by Rabault et al.Rabault et al. (2020), provides the main structure, and further details can be explored in their work. Figure 4 illustrates the fundamental framework of DRL, including the computational process of a single training episode and the utilized formulas. We refer to the work of Rabault et al.Rabault et al. (2020) and use a multilayer perceptron with two fully connected layers of 512 neurons each as the policy network. This network structure is sufficient to handle the complexity of the flow control problem in this study. Simpler network structures lack the modeling capacity to capture the mapping relationships, while more complex network structures are harder to train. The architecture of the ANN needs to be designed based on the dimensions of the input and output data, which often relies on empirical knowledge.
@@
The control action $a_{t}$ involves blowing and suction of two synthetic jets on the elliptical cylinder. The flow rates of the upper and lower synthetic jets are constrained such that their sum is zero, i.e., $Q_{\text{lower}}+Q_{\text{upper}}=0$. They also conducted an evaluation of the momentum injected into the flow field by the four jets. The RL agent controls the velocity of the upper synthetic jet, while the lower synthetic jet has the same velocity but in the opposite direction. To ensure energy efficiency in AFC, the velocity of the synthetic jets is limited to less than 2% of the inlet mean velocity, preventing the agent from utilizing excessively high jet velocities for drag reduction. Each control action, known as a ‘viscous action’, has a duration of 50 time steps. To ensure smooth transitions between consecutive actions, a smoothing function $S$ is introduced:
@@
where $C_{D,0}$ represents the baseline drag coefficient, serving as a reference point for drag reduction. $(C_{D})_{T_{i}}$ denotes the drag coefficient at time step $T_{i}$, with the objective of minimizing this value relative to the baseline. $(C_{L})_{T_{i}}$ signifies the lift coefficient at time step $T_{i}$, whose absolute value is penalized to mitigate lift forces that may destabilize the flow around the bluff body. $\omega$ is a weighting factor that quantifies the trade-off between minimizing drag and controlling lift fluctuations. Rabault and KuhnleRabault and Kuhnle (2019a) found that in the absence of a lift penalty, DRL produced a "cheating" strategy that resulted in reduced drag but was accompanied by a significant increase in lift. By incorporating a lift penalty in the reward function, this issue can be effectively mitigated. In previous work, the lift coefficient penalty factor is often set to 0.1Wang and Xu (2024a), 0.2Rabault et al. (2019); Tang et al. (2020) or 1Ren, Rabault, and Tang (2021). In this study, we use a penalty factor of 0.2, consistent with Rabault et al.Rabault et al. (2019)
The time granularity parameter significantly affects the agent’s performance and the complexity of the learning task. Precise adjustment to this hyperparameter, based on prior research and ongoing strategy refinement, is essential for achieving optimized control outcomes. In particular, each action in our analysis is applied for a duration of 0.025 seconds, which corresponds to 25 numerical simulation time steps. Considering that the Strouhal number for the flow around the elliptical cylinder in this study is between 0.3 and 0.4, we refer to the work of Rabault and KuhnleRabault and Kuhnle (2019a), where the duration of an episode corresponds to approximately 8 vortex shedding periods. Additionally, the work of Rabault et al.Rabault et al. (2019) considers an episode duration equivalent to about 6.5 vortex shedding periods. Therefore, in this study, the duration of an episode for the DRL training for AFC on the elliptical cylinder corresponds to approximately 6.25 to 8.33 vortex shedding periods. Each episode consists of 100 control actions, resulting in a maximum total time for an episode of $T_{\max}=T_{100}=2.5$ non-dimensional time units. The length of episodes is intentionally chosen to encompass multiple shedding cycles, which facilitates effective learning of the control algorithm by allowing the agent to observe and adapt to the flow dynamics over several cycles. Rabault and KuhnleRabault and Kuhnle (2019a) studied the impact of action update frequency on control strategies and found that approximately $T_{a}/T_{K}=10\%$, where $T_{K}$ represents the vortex shedding frequency and $T_{a}$ represents the duration of the action. In this study, the ratio of action duration to vortex shedding frequency is around 6% to 8%. This ensures that the action update frequency is sufficiently high for the ANN to control the system effectively while allowing the system enough time to respond to the control actions.
@@
The DRL training and numerical simulation are conducted on a high-performance computing (HPC) system equipped with an Intel® Xeon® Platinum 8358 CPU, operating at 2.60 GHz, and comprising a total of 64 cores evenly distributed across two sockets, each containing 32 cores. Specifically, the numerical simulation is performed using version 8 of the open-source software platform OpenFOAM® version 8 Jasak et al. (2007). The deployment of DRL algorithms is based on theTensorforce library Schaarschmidt et al. (2018), while the definitions of artificial neural networks and gradient descent algorithms are facilitated using the TensorFlow open-source library Abadi et al. (2016). For the DRL implementation, the agent is loaded using the Tensorforce platform with the PPO algorithm, and the environment is constructed using the Gym interface Brockman et al. (2016) . Fig. 4 illustrates the use of CFD calculations as the environment for DRL training and shows the effects of using 60 parallel CFD environments to accelerate the training process. This parallel strategy is inspired by the studies of Rabault and KuhnleRabault and Kuhnle (2019a) and Wang et al.Wang et al. (2022a), and also incorporateds insights from the work of Wang et al.Wang et al. (2022a) For a more comprehensive understanding, readers are encouraged to consult their works and the references therein.